OpenStudy (anonymous):

(8th grade Math) I dont understand it!! What is the slope of a line perpendicular to the line with equation y = 4x + 5? a.4 b.-4 c.-5 d.-1/4

OpenStudy (anonymous):

perpendicular means 90 degrees to each other the rule is \[gradient \times gradient_{perpendicular}=-1\] \[m \times m_{p}=-1\]

OpenStudy (anonymous):

\[4 \times m_p=-1\]